# Milky Way Globular Clusters: Nurseries for Dynamically-Formed Binary Black Holes

### v1

@@ 2.1 Galaxy Formation Simulation: GAMESH
The galaxy formation model GAMESH (Graziani+2015; Graziani+2017; Graziani+2020) combines an N-body simulation – which traces the assembly of dark matter – with a semi-analytic model that predicts star formation and the evolution of baryonic matter through processes such as chemical enrichment and radiative transfer. [^1] It is worth noting that, to enable a direct comparison with the results of (Graziani+2020), we adopt the same galaxy formation run, employing a semi-analytic prescription for radiative feedback, rather than executing the numerical radiative transfer component of the pipeline. Our semi-numerical approach focuses on a cube-shaped region measuring 4 comoving Mpc (cMpc) per side, centered on a halo analogous to the Milky Way (MW). This volume, designed to resemble the LG, is sufficiently large to encompass a statistically significant population of dwarf galaxies, enabling us to test our semi-analytic prescriptions for GC formation across a diverse range of astrophysical environments. The simulation achieves a mass resolution of approximately $3.4\,\times\,10^{5}$ M, allowing us to track the dynamics of mini-halos composed of few tens of dark matter particles. The N-body simulation tracks the formation of cosmic structures from redshift $z=20$ down to the present day ($z=0$), with snapshots saved at intervals of 15 Myr for $z\geq 10$, and every 100 Myr for $z<10$. In each snapshot, dark matter halos are identified using a Friends-of-Friends algorithm with a linking length of 0.2 and a minimum of 100 particles per halo. A particle-based merger tree is then constructed to trace the hierarchical assembly of these halos over cosmic time. Because our simulation focuses on a volume representative of the LG, the rates of star and cluster formation within it tend to be higher than the cosmic average. Consequently, when comparing our results to studies based on larger, more representative cosmological volumes (e.g., Hopkins+2014; McAlpine+2016; Kaviraj+2017; Donnari+2019; Venditti+2023), our predictions for certain quantities should be interpreted as upper-limit estimates (Mapelli+2018; Artale+2019; Broekgaarden+2022; Franciolini+2024; Bruel+2024; Liu+2024). Finally, a key advantage of testing our novel coupling approach within a MW-like galaxy is the exceptional availability of observational data. The existence of comprehensive GC catalogs for the MW enables detailed and rigorous comparisons with our simulation outcomes (Forbes+2010; Dotter+2010; Dotter+2011; VandenBerg+2013), providing a robust framework for validating the model’s ability to reproduce the fundamental properties of galactic GCs.
@@ 2.2 Coupling setup
    Zoom-in simulations (Ma+2020; Chen+2021; Fukushima+2021) suggest that GC formation is most likely to occur when the gas surface density of giant molecular clouds exceeds approximately $10^{3}\,M_{\odot}\,{\rm pc}^{-2}$ (see Adamo+2020; Kruijssen+2025, for reviews). Within the GAMESH framework, we estimate the gas surface density by dividing the total gas mass $M_{\rm g}$ contained in a dark matter halo by the effective area of the galaxy:
    $\Sigma_{\rm g}=\frac{M_{\rm g}}{\pi\,r^{2}_{\rm h}}>10^{3}\,M_{\odot}\,{\rm pc}^{-2},$ (1)
    where $r_{\rm h}$ is the half-mass stellar radius. Although we apply this gas surface density threshold on galactic scales, the approximation is particularly valid at high redshifts, where $r_{\rm h}$ is comparable to the typical size of giant molecular clouds ($\sim 10\div 200$ pc, Dame+2001). At lower redshifts, where these scales diverge more significantly, the impact on our results remains limited, as the bulk of GC formation occurs at high redshifts (z ¿ 3). To compute $r_{\rm h}$ from the stellar mass $M_{\star}$ of each dark matter halo, we use the mass–size relation from (Shen+2003), incorporating its redshift evolution as described in (vanDerWel+2014).
@@
    For each galaxy which can host GC formation, we compute the cluster formation efficiency $\Gamma$ (Li+2017; Pfeffer+2019; Lahen+2020; Li+2022; Grudic+2023) following the work of (Kruijssen+2012):
    $\Gamma=\left(1.15+0.6\,\Sigma_{\rm SFR}^{-0.4}+0.05\,\Sigma^{-1}_{\rm SFR}\right)^{-1}$ (2)
    where $\Sigma_{\rm SFR}=\frac{\rm SFR}{\pi\,r_{h}^{2}}$ is the star formation rate surface density. A fraction of the stellar mass that forms between two snapshots, $\Delta M_{\star}={\rm SFR}\,\Delta t$, is assumed to be bound in GCs ($M_{\rm cl}$):
    $M_{\rm cl}=\Gamma\,\Delta M_{\star}\,.$ (3)
@@
    We sample GC masses from a cluster initial mass function modeled as a power law with an exponent of 2-2 (Li+2017; Li+2018; Pfeffer+2019; Li+2022; Lahen+2024). Since the CPS prescriptions are typically less accurate for lighter clusters, the minimum GC mass is commonly set at $m_{\rm cl,min}=10^{5}\,M_{\odot}$. However, for the scope of this study, we extend this minimum mass to $10^{4}\,M_{\odot}$, based on the assumption that clusters with masses below $10^{5}\,M_{\odot}$ do not contribute significantly to the overall BBH merger rate. The maximum mass, conversely, depends on the properties of the host galaxy, approximately scaling as (Li+2022):
    $m_{\rm cl,max}=\Sigma_{\rm SFR}^{2/3},$
    We randomly sample GC masses from the cluster initial mass function until the remaining stellar mass bound in GC falls below $m_{\rm cl,min}$.
    The initial metallicity of GCs is assumed to be the same as that of their parent galaxy. Their initial half-mass radii are drawn from a uniform distribution ranging between 0.1 pc and 1 pc Askar+2017; Kremer+2020a; ArcaSedda+2020. Although these initial conditions are informed by both simulation results and observational data, a significant uncertainty remains regarding the binary fraction within GCs (Richard+2015; Stanway+2020). In this study, we adopt a binary fraction of unity for all clusters and leave a detailed exploration of its effects for future work.
@@
    Once formed within galactic disks, GCs are not immune to disruption. They are frequently subjected to strong tidal forces from dense gas clouds, as demonstrated in various studies (Gieles+2006; Elmegreen+2010; Kruijssen+2011). Following the framework presented by (Kruijssen+2015), we model this rapid disruption process – referred to as the “cruel cradle effect” – as:
@@
    where $t_{\mathrm{cce}}$ denotes the characteristic timescale of the cruel cradle effect (Kruijssen+2012b). This timescale depends on various galactic properties and is expressed as (Kruijssen+2015):
    $t_{\mathrm{cce}}=176\,\left(\frac{f_{\Sigma}}{4}\right)^{-1}\left(\frac{\rho_{\mathrm{ISM}}}{M_{\odot}\,\mathrm{pc}^{-3}}\right)^{-3/2}\left(\frac{m_{\rm cl}}{10^{5}\,M_{\odot}}\right)\,\phi_{\mathrm{ad}}^{-1}\,\mathrm{Myr},$ (5)
@@
    • $\rho_{\mathrm{ISM}}=\frac{3\,\Omega^{2}(r_{\rm g})}{\pi\,G}$ is the average mid-plane density of the interstellar medium in a stable galactic disk (Krumholz+2005; DeLucia+2024), evaluated at the galactocentric radius $r_{\rm g}$ of the GC. In this study, we assume GCs to be on circular orbits and located at a radius $r_{\rm g}=r_{\rm h}$;
@@
    Finally, we apply a GC survival prescription during galaxy mergers, following the models presented in (Kruijssen+2011; Kruijssen+2015; DeLucia+2024). After the coalescence of two galaxies, we compute the fraction $f_{\mathrm{surv}}$ of GCs that migrate from the galactic disks of their progenitor galaxies - where they typically form - to the halo of the merger remnant, as described in (Kruijssen+2012b):
    $f_{\mathrm{surv}}=4.5\times 10^{-8}\left(\frac{m_{\mathrm{cl,min}}}{100\,M_{\odot}}\right)^{2}\left(\frac{t_{\mathrm{depl}}}{\mathrm{yr}}\right)^{0.77-0.22\log(m_{\mathrm{cl,min}}/100\,M_{\odot})},$ (6)
    where $t_{\mathrm{depl}}$ is the gas depletion timescale, estimated as the ratio of the gas mass to the peak star formation rate (SFR) during the merger-induced starburst (Kruijssen+2012b). Following (DeLucia+2024), we approximate this peak SFR by summing the SFRs of the two merging galaxies at the snapshot in which the merger is identified.
    Once relocated to the halo, GCs are assumed to evolve in isolation and are unaffected by any subsequent mergers (Lamers+2010). The migrating GCs are randomly selected from those residing in the progenitor galaxies’ disks prior to the merger.
@@
    Incorporating these prescriptions is essential to avoid overestimating the number of GCs in the MW halo that would survive to the present day.
@@ 2.3 Cluster Population Synthesis Codes: RAPSTER and FASTCLUSTER
To generate large populations of GCs and study their dynamically-formed BBHs, the only practical approach both in terms of computational efficiency and statistical robustness is to employ CPS codes such as RAPSTER (Kritos+2023; Kritos+2024) and FASTCLUSTER (Mapelli+2021; Torniamenti+2024). These tools adopt a semi-analytic framework specifically designed to model the coupled evolution of massive star clusters ($m_{\mathrm{cl}}\geq 10^{5}\,M_{\odot}$) and their resident BBH populations. The main advantage lies in the ability to simulate a large number of clusters efficiently, a critical feature for statistical studies of dynamically-formed BBHs within cosmological simulations.
@@
The codes are built upon a set of physically motivated prescriptions calibrated against direct N-body and Monte Carlo simulations. These describe all relevant dynamical channels, such as three-body interactions, GW captures, and binary-single scatterings. Their simplified prescriptions circumvent the computational intensity of direct $N$-body simulations, allowing to evolve individual GCs with $m_{\mathrm{cl}}\geq 10^{5}\,M_{\odot}$ within a matter of minutes. However, the computational speed comes at a cost: only the average properties of the cluster and its BHs are evolved with time, whereas the internal dynamics of the cluster is not integrated directly.
@@
Their considerable flexibility enables to explore a wide range of input parameters, including BBH masses, spins, initial orbital eccentricities, delay times to merger, and various properties of the host star clusters, such as their masses, densities, and initial binary fractions. Furthermore, these CPS codes allow for the temporal evolution of the host stellar cluster itself, incorporating key astrophysical processes such as stellar mass loss, two-body relaxation, and tidal stripping by the host galaxy. While both codes are effective CPS tools, their main differences lie in their primary focus. With its fast Monte-Carlo sampling algorithm and extensive calibration on GC data, FASTCLUSTER aims to reproduce statistically time-dependent observations like the merger rate of BBHs when coupled with analytic star formation histories. Conversely, RAPSTER is a more unconstrained model with several free parameters, initially developed to investigate the fundamental possibility of producing massive and intermediate-mass BHs from globular and nuclear star clusters. Despite these differences, both codes are specifically optimized for GW astronomy, with the shared goal of creating extensive catalogs of dynamically-formed BBH mergers for comparison with observational data. For further information on how the implementation of this physics takes place, we refer to the main papers for each code. To compute the initial single-BH mass spectrum, both codes adopt the stellar evolution prescriptions from SEVN (Spera+2017; Iorio+2023), which determine the remnant mass as a function of the zero-age main-sequence mass of stars. This study focuses exclusively on GCs composed of Population II and I stars; thus, we sample initial stellar masses from a Kroupa initial mass function (IMF) (Kroupa+2001), spanning the range $0.08,\,200]\,M_{\odot}$ and covering metallicities in the interval $10^{-2},1]\,Z_{\odot}$. Given that our LG environment is overdense, metal enrichment in low-metallicity halos occurs extremely rapidly. As a consequence, only a few halos reach the conditions necessary for Population III GC formation. Their contribution to the overall results is negligible and, hence, not considered here; we defer their treatment to future work focused on high-redshift regimes.
@@
Although we adopt the same initial conditions for GCs, intrinsic differences between the working principle of CPS codes persist. Preliminary tests performed with non-calibrated code settings revealed significant structural differences in the algorithms, leading to substantial divergences in final outcomes. Specifically, variations in setting parameters like the initial binary fraction and the number of BH merger chains within each GC were identified as major sources of discrepancy.
@@
To address these code-specific structural differences and avoid highly divergent results, we implement a calibration method focusing on the most critical diverging element, i.e. the BH merger chains. More precisely, FASTCLUSTER requires the number of BH merger chains within each GC as an explicit input parameter, a value that RAPSTER does not explicitly demand but intrinsically calculates. We therefore extract this value from the RAPSTER outcomes and adopt the identical number in FASTCLUSTER for an accurate one-to-one comparison.

### v2

@@ 2.1 Galaxy formation simulation: GAMESH
The galaxy formation model GAMESH (Graziani et al. 2015, 2017, 2020c) combines an N-body simulation – which traces the assembly of dark matter – with a semi-analytic model that predicts star formation and the evolution of baryonic matter through processes such as chemical enrichment and radiative transfer. [^1] It is worth noting that, to enable a direct comparison with the results of Graziani et al. (2020c), we adopt the same galaxy formation run, employing a semi-analytic prescription for radiative feedback, rather than executing the numerical radiative transfer component of the pipeline. Our semi-numerical approach focuses on a cube-shaped region measuring 4 comoving Mpc (cMpc) per side, centered on a halo analogous to the Milky Way (MW). This volume, designed to resemble the LG, is sufficiently large to encompass a statistically significant population of dwarf galaxies, enabling us to test our semi-analytic prescriptions for GC formation across a diverse range of astrophysical environments. The simulation achieves a mass resolution of approximately $3.4\,\times\,10^{5}\,\rm M_{\odot}$, allowing us to track the dynamics of mini-halos composed of few tens of dark matter particles. The N-body simulation tracks the formation of cosmic structures from redshift $z=20$ down to the present day ($z=0$), with snapshots saved at intervals of $15\,\rm Myr$ for $z\geq 10$, and every $100\,\rm Myr$ for $z<10$. In each snapshot, dark matter halos are identified using a Friends-of-Friends algorithm with a linking length of 0.2 and a minimum of 100 particles per halo. A particle-based merger tree is then constructed to trace the hierarchical assembly of these halos over cosmic time. GAMESH incorporates several key baryonic processes. The mass of gas within each halo is initially assumed to be proportional to its dark matter mass, scaled by the cosmic baryon fraction. As halos evolve along the merger tree, they acquire gas through both smooth accretion from the surrounding environment and mergers with other structures. Consequently, the gas accretion rate is determined by the growth of the dark matter halo: as a halo increases its mass, it self-consistently acquires a proportional amount of cosmic gas (Graziani et al. 2015).
In each halo, the star formation rate (SFR) is computed as SFR = $\epsilon_{\star}$ $\frac{\rm M_{gas}}{\rm t_{dyn}}$, depending on the star formation efficiency $\epsilon_{\star}$, the total gas mass $\rm M_{gas}$ and the dynamical time of the host halo $\rm t_{dyn}\sim(1+z)^{-\frac{3}{2}}$ (Graziani et al. 2017). We adopt a star formation efficiency of $\epsilon_{\star}=0.02$, a value that has been calibrated within the GAMESH framework to match the observed properties of the LG. As shown in Fig. 1 of Graziani et al. (2017), this efficiency allows the model to successfully reproduce the cosmic SFR density evolution, ensuring that the predicted growth of stellar mass is consistent with observations at z = 0.
Mechanical feedback is regulated by supernova explosions, which drive galactic winds capable of ejecting gas into the intergalactic medium. The mass of the ejected gas is calculated as $\Delta\rm M=2\,\epsilon_{W}\,\frac{E_{SN}}{v_{esc}^{2}}$, where $E_{SN}$ is the energy released by supernovae, $\rm v_{esc}$ is the escape velocity of the halo, and $\epsilon_{W}$ represents the wind efficiency, which controls how effectively supernova energy is converted into kinetic energy to drive outflows (Graziani et al. 2017).
We model the chemical enrichment of the interstellar medium by implementing a simplified supernova feedback prescription, following Salvadori et al. (2008). In line with the instantaneous recycling approximation (Tinsley 1980), we assume that metals and gas are mixed both immediately and uniformly (Salvadori et al. 2007, 2010).
Radiative feedback is treated following the simplified approach in Graziani et al. (2017), by assuming an instantaneous reionization redshift at z = 6. Because our simulation focuses on a volume representative of the LG, the rates of star and cluster formation within it tend to be higher than the cosmic average. Consequently, when comparing our results to studies based on larger, more representative cosmological volumes (e.g., Hopkins et al. 2014; McAlpine et al. 2016; Kaviraj et al. 2017; Donnari et al. 2019; Venditti et al. 2023), our predictions for certain quantities should be interpreted as upper-limit estimates (Mapelli et al. 2018; Artale et al. 2019; Broekgaarden et al. 2022; Franciolini et al. 2024; Bruel et al. 2024; Liu et al. 2024). Finally, a key advantage of testing our novel coupling approach within a MW-like galaxy is the exceptional availability of observational data. The existence of comprehensive GC catalogs for the MW enables detailed and rigorous comparisons with our simulation outcomes (Forbes and Bridges 2010; Dotter et al. 2010, 2011; VandenBerg et al. 2013), providing a robust framework for validating the model’s ability to reproduce the fundamental properties of galactic GCs.
@@ 2.2 Coupling setup
    Zoom-in simulations (Ma et al. 2020; Chen et al. 2021; Fukushima and Yajima 2021) suggest that GC formation is most likely to occur when the gas surface density of giant molecular clouds exceeds approximately $10^{3}\,\rm M_{\odot}\,\rm pc^{-2}$ (see Adamo et al. 2020; Kruijssen 2026, for reviews). Recent JWST observations of lensed galaxies have provided empirical support for this scenario, revealing the presence of ultra-compact proto-globular clusters with stellar surface densities reaching $\sim 10^{5}\,\rm M_{\odot}\,\rm pc^{-2}$ (Adamo et al. 2024). Within the GAMESH framework, we estimate the gas surface density by dividing the total gas mass $\rm M_{\rm g}$ contained in a dark matter halo by the effective area of the galaxy
    $\Sigma_{\rm g}=\frac{\rm M_{\rm g}}{\pi\,\rm r^{2}_{\rm h}}>10^{3}\,\rm M_{\odot}\,{\rm pc}^{-2},$ (1)
    where $\rm r_{\rm h}$ is the half-mass stellar radius. The exact gas surface density threshold for GC formation remains a subject of debate. Hosting galaxies are expected to be highly fragmented (Fujimoto et al. 2025), implying that local density peaks can trigger GC formation even when the disk-averaged density remains below the threshold. Although we apply this gas surface density threshold on galactic scales, the approximation is particularly valid at high redshifts, where $\rm r_{\rm h}$ is comparable to the typical size of giant molecular clouds ($\sim 10-200$ pc, Dame et al. 2001). At lower redshifts, where these scales diverge more significantly, the impact on our results remains limited, as the bulk of GC formation occurs at high redshifts (z ¿ 3). It is important to stress that recent high-resolution JWST observations of lensed high-redshift galaxies have discovered several bound clusters within regions smaller than 100 pc (Bradač et al. 2025; Messa et al. 2025), confirming that at cosmic dawn the GC progenitors form at scales comparable to the typical size of giant molecular clouds.
    To compute $r_{\rm h}$ from the stellar mass $M_{\star}$ of each dark matter halo, we use the mass–size relation from Shen et al. (2003), incorporating its redshift evolution as described in van der Wel et al. (2014). This scaling is consistent with recent JWST observations confirming that galaxies at z >> 6 are significantly more compact than their local counterparts (Mowla et al. 2024; Kalita et al. 2025).
    We have performed a sensitivity test on the choice of this threshold and the results are discussed in Appendix C.
@@
    For each galaxy that can host GC formation, we compute the cluster formation efficiency $\Gamma$ (Li et al. 2017; Pfeffer et al. 2019; Lahén et al. 2020; Li et al. 2022; Grudić et al. 2023) following the work of Kruijssen (2012):
    $\Gamma=\left(1.15+0.6\,\Sigma_{\rm SFR}^{-0.4}+0.05\,\Sigma^{-1}_{\rm SFR}\right)^{-1},$ (2)
    where $\Sigma_{\rm SFR}=\frac{\rm SFR}{\pi\,\rm r_{h}^{2}}$ is the star formation rate surface density. A fraction of the stellar mass that forms between two snapshots, $\Delta\rm M_{\star}={\rm SFR}\,\Delta t$, is assumed to be bound in GCs ($\rm M_{\rm cl}$):
    $\rm M_{\rm cl}=\Gamma\,\Delta\rm M_{\star}\,.$ (3)
@@
    We sample GC masses from a cluster initial mass function modeled as a power law with an exponent of 2-2 (Li et al. 2017, 2018; Pfeffer et al. 2019; Li et al. 2022; Lahén et al. 2024). This choice is further supported by recent JWST observations of star clusters across cosmic time (z $\approx$ 1-9), which find a mass function slope of $1.89^{+0.13}_{-0.12}$, consistent with a universal power-law index of 2-2 (Claeyssens et al. 2026).
    Since the CPS prescriptions are typically less accurate for lighter clusters, the minimum GC mass is commonly set at $m_{\rm cl,min}=10^{5}\,\rm M_{\odot}$. However, for the scope of this study, we extend this minimum mass to $10^{4}\,\rm M_{\odot}$, based on the assumption that clusters with masses below $10^{5}\,\rm M_{\odot}$ do not contribute significantly to the overall BBH merger rate. The maximum mass, conversely, depends on the properties of the host galaxy, approximately scaling as (Li et al. 2022):
    $m_{\rm cl,max}=\Sigma_{\rm SFR}^{2/3}.$
    We randomly sample GC masses from the cluster initial mass function until the remaining stellar mass bound in GC falls below $m_{\rm cl,min}$. The initial metallicity of GCs is assumed to be the same of their parent galaxy. Their initial half-mass radii are drawn from a uniform distribution ranging between 0.1 pc and 1 pc (Askar et al. 2017; Kremer et al. 2020b; Arca Sedda et al. 2020). Although these initial conditions are informed by both simulation results and observational data, a significant uncertainty remains regarding the binary fraction within GCs (de Grijs et al. 2015; Stanway et al. 2020). In this study, we adopt a binary fraction of unity for all clusters and leave a detailed exploration of its effects for future work.
@@
    Once formed within galactic disks, GCs are not immune to disruption. They are frequently subjected to strong tidal forces from dense gas clouds, as demonstrated in various studies (Gieles et al. 2006; Elmegreen and Hunter 2010; Kruijssen et al. 2011). Following the framework presented by Kruijssen (2015), we model this rapid disruption process – referred to as the ”cruel cradle effect” – as:
@@
    where $t_{\mathrm{cce}}$ denotes the characteristic timescale of the cruel cradle effect (Kruijssen et al. 2012). This timescale depends on various galactic properties and is expressed as (Kruijssen 2015):
    $t_{\mathrm{cce}}=176\,\left(\frac{f_{\Sigma}}{4}\right)^{-1}\left(\frac{\rho_{\mathrm{ISM}}}{\rm M_{\odot}\,\mathrm{pc}^{-3}}\right)^{-3/2}\left(\frac{m_{\rm cl}}{10^{5}\,\rm M_{\odot}}\right)\,\phi_{\mathrm{ad}}^{-1}\,\mathrm{Myr},$ (5)
@@
    • $\rho_{\mathrm{ISM}}=\frac{3\,\Omega^{2}(r_{\rm g})}{\pi\,G}$ is the average mid-plane density of the interstellar medium in a stable galactic disk (Krumholz and McKee 2005; De Lucia et al. 2024), evaluated at the galactocentric radius $r_{\rm g}$ of the GC. In this study, we assume GCs to be on circular orbits and located at a radius $r_{\rm g}=r_{\rm h}$. Although this represents a simplified orbital model, we perform a dedicated sensitivity analysis in Appendix B to assess the impact of this assumption on our results and to quantify any potential systematic effects;
@@
    Finally, we apply a GC survival prescription during galaxy mergers, following the models presented in Kruijssen et al. (2011); Kruijssen (2015); De Lucia et al. (2024). After the coalescence of two galaxies, we compute the fraction $f_{\mathrm{surv}}$ of GCs that migrate from the galactic disks of their progenitor galaxies - where they typically form - to the halo of the merger remnant, as described in Kruijssen et al. (2012):
    $f_{\mathrm{surv}}=4.5\times 10^{-8}\left(\frac{m_{\mathrm{cl,min}}}{100\,\rm M_{\odot}}\right)^{2}\left(\frac{t_{\mathrm{depl}}}{\mathrm{yr}}\right)^{0.77-0.22\log(m_{\mathrm{cl,min}}/100\,\rm M_{\odot})},$ (6)
    where $t_{\mathrm{depl}}$ is the gas depletion timescale, estimated as the ratio of the gas mass to the peak star formation rate (SFR) during the merger-induced starburst (Kruijssen et al. 2012). Following De Lucia et al. (2024), we approximate this peak SFR by summing the SFRs of the two merging galaxies at the snapshot in which the merger is identified.
    Once relocated to the halo, GCs are assumed to evolve in isolation and are unaffected by any subsequent mergers (Lamers et al. 2010). The migrating GCs are randomly selected from those residing in the progenitor galaxies’ disks prior to the merger.
@@
    We clarify that the disruption mechanisms described by Equations 4 and 6 are applied to the initial cluster population, before their evolution is tracked via the CPS models. Specifically, this preliminary analysis allows us to determine the lifetimes of GCs. Once this preliminary selection is made, the subsequent mass evolution tracked within the CPS models focuses exclusively on the mass loss induced by the external galactic potential.
@@ 2.3 Cluster population synthesis codes: RAPSTER and FASTCLUSTER
To generate large populations of GCs and study their dynamically formed BBHs, the only practical approach both in terms of computational efficiency and statistical robustness is to employ CPS codes such as RAPSTER (Kritos et al. 2023b, 2024) and FASTCLUSTER (Mapelli et al. 2021; Torniamenti et al. 2024). These tools adopt a semi-analytic framework specifically designed to model the coupled evolution of massive star clusters ($m_{\mathrm{cl}}\geq 10^{5}\,\rm M_{\odot}$) and their resident BBH populations. The main advantage lies in the ability to simulate a large number of clusters efficiently, a critical feature for statistical studies of dynamically formed BBHs within cosmological simulations.
@@
The codes are built upon a set of physically motivated prescriptions calibrated against direct N-body and Monte Carlo simulations. These describe all relevant dynamical channels, such as three-body interactions, GW captures, and binary-single scatterings. Their simplified prescriptions circumvent the computational intensity of direct $N$-body simulations, allowing to evolve individual GCs with $m_{\mathrm{cl}}\geq 10^{5}\,\rm M_{\odot}$ within a matter of minutes. However, the computational speed comes at a cost: only the average properties of the cluster and its BHs are evolved with time, whereas the internal dynamics of the cluster is not integrated directly. By employing both RAPSTER and FASTCLUSTER, we aim to quantify the impact of different semi-analytical prescriptions on our results, ensuring that our findings regarding dynamically formed BBHs remain robust across different modeling approaches within the same cosmological framework.
@@
Their considerable flexibility enables to explore a wide range of input parameters, including BBH masses, spins, initial orbital eccentricities, delay times to merger, and various properties of the host star clusters, such as their masses, densities, and initial binary fractions. Furthermore, these CPS codes allow for the temporal evolution of the host stellar cluster itself, incorporating key astrophysical processes such as stellar mass loss, two-body relaxation, and tidal stripping by the host galaxy. While both codes are effective CPS tools, their main differences lie in their primary focus. With its fast Monte-Carlo sampling algorithm and extensive calibration on GC data, FASTCLUSTER aims to reproduce statistically time-dependent observations, such as the merger rate of BBHs when coupled with analytic star formation histories. Conversely, RAPSTER is a more unconstrained model with several free parameters, initially developed to investigate the fundamental possibility of producing massive and intermediate-mass BHs from globular and nuclear star clusters. Despite these differences, both codes are specifically optimized for GW astronomy, with the shared goal of creating extensive catalogs of dynamically formed BBH mergers for comparison with observational data. For further information on how the implementation of this physics takes place, we refer to the main papers for each code. To compute the initial single-BH mass spectrum, both codes adopt the stellar evolution prescriptions from SEVN (Spera and Mapelli 2017; Iorio et al. 2023), which determine the remnant mass as a function of the zero-age main-sequence mass of stars. This study focuses exclusively on GCs composed of Population II and I stars; thus, we sample initial stellar masses from a Kroupa initial mass function (IMF) (Kroupa 2001), spanning the range $0.08,\,200]\,\rm M_{\odot}$ and covering metallicities in the interval $10^{-2},1]\,Z_{\odot}$. Given that our LG environment is overdense, metal enrichment in low-metallicity halos occurs extremely rapidly. As a consequence, only a few halos reach the conditions necessary for Population III GC formation. Their contribution to the overall results is negligible and, hence, not considered here; we defer their treatment to future work focused on high-redshift regimes.
@@
Although we adopt the same initial conditions for GCs, intrinsic differences between the working principle of CPS codes persist. Preliminary tests performed with non-calibrated code settings revealed significant structural differences in the algorithms, leading to substantial divergences in final outcomes. Specifically, variations in setting parameters such as the initial binary fraction and the number of BH merger chains within each GC were identified as major sources of discrepancy.
@@
To address these code-specific structural differences and avoid highly divergent results, we implement a calibration method focusing on the most critical diverging element, i.e., the BH merger chains. More precisely, FASTCLUSTER requires the number of BH merger chains within each GC as an explicit input parameter, a value that RAPSTER does not explicitly demand but intrinsically calculates. We therefore extract this value from the RAPSTER outcomes and adopt the identical number in FASTCLUSTER for an accurate one-to-one comparison.